# A robust cosmic standard ruler from the cross-correlations of galaxies and dark sirens

### v1

@@ Abstract
Observations of gravitational waves (GWs) from dark sirens allow us to infer their locations and distances. Galaxies, on the other hand, have precise angular positions but no direct measurement of their distances – only redshifts. The cross-correlation of GWs, which we limit here to binary black hole mergers (BBH), in spherical shells of luminosity distance $D_{L}$, with galaxies in shells of redshift $z$, leads to a direct measurement of the Hubble diagram $D_{L}(z)$. Since this standard ruler relies only on the statistical proximity of the dark sirens and galaxies (a general property of large-scale structures), it is essentially model-independent: the correlation is maximal when both redshift and $D_{L}$ shells coincide. We forecast the constraining power of this technique, which we call Peak Sirens, for run 5 (O5) of LIGO-Virgo-KAGRA (LVK), as well as for the third-generation experiments Einstein Telescope and Cosmic Explorer. We employ thousands of full-sky light cone simulations with realistic numbers for the tracers, and include masking by the Milky Way, lensing and inhomogeneous GW sky coverage. We find that the method is not expected to suffer from some of the issues present in other dark siren methods, such as biased constraints due to incompleteness of galaxy catalogs or dependence on priors for the merger rates of BBH. We show that with Peak Sirens, given the projected O5 sensitivity, LVK can measure $H_{0}$ with $7\%$ precision by itself, assuming $\Lambda$CDM, and $4\%$ precision using external datasets to constrain $\Omega_{m}$. We also show that future third-generation GW detectors can achieve, without external data, sub-percent uncertainties in $H_{0}$ assuming $\Lambda$CDM, and 3% in a more flexible $w_{0}w_{a}$ model. The method also shows remarkable robustness against systematic effects such as the modeling of non-linear structure formation.
@@ 1 Introduction
The large-scale structures of the Universe are increasingly being mapped with multiple tracers and messengers, and the complementarity between different surveys has enormous potential to test cosmological models and resolve some of the existing tensions of the standard $\Lambda$CDM scenario [1, 2]. In this paper we investigate how we can exploit the cross-correlations between redshift surveys and detections of gravitational waves (GW) from dark sirens (most of which are binary black hole mergers, BBH) to trace the Hubble distance diagram $D_{L}(z)$ – even when the galaxies that host GWs are absent from the redshift survey. The idea is illustrated in Fig. 1: a galaxy map, where the radial information is only in the redshifts, covers the same area in the sky as a map of GWs, for which the gravitational wave (luminosity) distances can be inferred from the waveforms.
@@ 2.1 Multi-tracer likelihoods
2.2 A fast estimator for redshift-distance cross-correlations
@@
$a_{\ell m}^{b}(D_{i},\theta)\,=\sum_{D_{\lambda}\in D_{i}}\,a_{\ell m}^{b}(D_{\lambda},\theta)\,.$ (2.12)
Suppose that the luminosity distance bins of our data set are fixed – that is, independent of the cosmology. In practice this is how we would proceed to compute the cross correlations of a set of observed BBHs with a set of galaxies in fixed redshift bins. Now let’s say that the small bins $D_{\lambda}$ are not fixed, but rather we fix some set of very thin redshift bins $z_{\lambda}$. As the connection between redshift and distance is mediated by the cosmological model, we can write $D_{\lambda}=D_{L}(z_{\lambda},\theta)\,$. Therefore, we can write the harmonic mode in a distance bin $D_{i}$ as a sum over the redshift bins $z_{\lambda}$ that fit into that distance bin, for the given cosmology:
$a_{\ell m}^{b}(D_{i},\theta)\,=\sum_{z_{\lambda}\in\,z(D_{i},\theta)}\,a_{\ell m}^{b}(z_{\lambda},\theta)\,.$ (2.13)
Now say that we change the set of cosmological parameters from ${\theta}_{\rm fid}$ – where fid represents a fiducial set of parameters – to another arbitrary set ${\theta}$. Although the amplitudes $a_{\ell m}^{b}(z_{\lambda},\theta)$ change, the correlation function still peaks when the redshift of the galaxies coincide with $z_{\lambda}$. The most salient effect of the change in the parameters is in fact a shift in the way that the very thin redshift bins ($z_{\lambda}$) are mapped into the distance bins ($D_{i}$), which change the interval of the sum over redshifts for any given distance bin in Eq. (2.13).
Let us then define a set of harmonic coefficients in the new cosmological model by:
$a_{\ell m}^{b}(D_{i},\theta)\,=\sum_{z_{\lambda}\in\,z(D_{i},\theta)}\,a_{\ell\,m}^{b}(z_{\lambda},\theta_{\rm fid})\,.$ (2.14)
@@
where $C^{gb}_{\ell}(z_{j},z_{\lambda},\theta_{\rm fid})$ is the cross-angular power spectrum computed in thin redshift bins $z_{\lambda}$ (both for galaxies and BBHs) and in a fiducial cosmology $\theta_{\rm fid}$, averaged over a given number of realizations (for our purposes, of the order of thousands, as we will see). The only general condition is that any distance bins are wide enough to comprise several thin redshift bins, so that our analysis is not significantly affected by that spatial resolution.
Equation (2.15) is the key approximation that makes our analysis possible. The main reason it works is because of the coincidence of the peak of the cross-correlation with the Hubble diagram, which is a generic feature of the large-scale structures and as such should not depend significantly on the galaxy or BBH populations, bias modeling or non-linearities. In essence, the log-likelihood (2.1) will be sensitive mostly to the peak of the cross-correlation signal, broadened by the uncertainty in the BBH localization, and this is overwhelmingly dictated by the distance-redshift relation. The validity of this approximation will be demonstrated in practice in Section 4.2.
@@ 3 Simulations
In order to assess the potential of the Peak Sirens method to constrain cosmology we forward-model the cosmological distribution of galaxies and BBHs. The starting point is a full-sky cosmological mock of the past light cone, possible thanks to the publicly available GLASS code [43]. We used GLASS to simulate 1000 lognormal galaxy mocks in very thin redshift bins, of width $dz=0.02$, over the interval $z\in[0,2]$ – resulting in the mean redshift bins $\bar{z}$:
@@ 3.1 Gravitational wave Fisher forecasts
We performed Fisher forecasts using the code GWDALI [63], with the following 11 free parameters:
@@
where $M_{c}$ is the redshifted chirp mass, $\eta$ is the symmetric mass ratio, and $t_{c}$ and $\phi_{c}$ are the time and phase of coalescence. We considered as threshold for detection having the total SNR of the network greater than 12. For the LVK+ET network, we have further imposed that at least one of the LVK detectors have SNR > 1. That does not mean, therefore, that the events included in this case would actually be detected by the 2nd generation network. Rather, LVK would help a 3rd generation detector (here, the triangular ET) to triangulate the signal improving localization knowledge, similarly to how Virgo has helped LIGO on localizing sources in their first three observing runs even if for the events of which Virgo alone can not claim a detection. After these cuts in SNR, we prospect the total amount of 377 oberved events for LVK O5, while for LVK+ET and ET+2CE we expect this number to grow to, respectively, 54991 and 73260 detected events.
@@
Figure 3: Cumulative distribution of forecast sky position uncertainties (left) and relative distance errors (right) for the three GW detector networks here considered. The number of events are much higher for LVK+ET compared to LVK O5, especially at high $z$, a fraction of which with poor localization, result in in a modest improvement in the average localization. See also Figure 4.
@@ 3.2 Galaxy and BBH mocks
with $z_{0}=0.7$, $\alpha=2.0$, $\beta=1.5$, and a linear galaxy bias, $b_{g}(z)=(1+z)$. With our choice of parameters (including the normalization of the Smail distribution) we obtained approximately 34 million galaxies up to $z=2$, a number that will easily be surpassed by upcoming surveys. To account for partial sky coverage, we apply a mask to the galaxy maps, restricting the analysis to latitudes $|\phi|\geq 19^{\circ}$, approximately $\nicefrac{{1}}{{3}}$ of the sky.
@@ 3.2.3 Weak lensing effects
Fig. 7 can also be directly compared with the analytical prediction of Ref. [4] – see Fig. 1 of that paper. Apart from an overall normalization, the results for the weak-lensing contribution on the signal of $C_{\ell}^{gb}$ for the optimal case of ET+2CE are consistent with that analytical estimate. However, due to the long tail on the convergence PDFs, we observe a non-negligible correlation between BBHs and galaxies at redshifts far below $z_{g}=1$. One of the main differences between our results and those of Ref. [4] is the steep drop of the lensing signal as $z\to 0$, which is due to the fact that the convergence PDFs become narrower and nearly Gaussian as we $z\to 0$, corresponding to a fast decreasing fraction of highly magnified objects.
@@ 4.1 Forecasts for cosmological parameters
The distance bins are kept fixed but can be defined arbitrarily, independently of the redshift bins. For this analysis we choose 50 distance bins ranging from $D_{\text{fid}}(z=0)$ to $D_{\text{fid}}(z=2)$, in equally spaced redshift intervals, where by $D_{\text{fid}}$ we mean the luminosity distance calculated in the fiducial cosmology. This binning strategy offers two key advantages: first, it ensures that each distance bin includes one or more thin redshift bins in the discrete sum of Eq. (2.15). Second, the cross-correlation in the fiducial cosmology becomes a diagonal stripe, which simplifies the computation of the inverse correlation matrices in Eq. (2.1).
@@
For middle- and long-term scenarios we investigate LVK+ET and ET+2CE as potential networks of GW detectors. For these forecasts we also consider, besides $\Lambda$CDM, a second model, $w_{0}w_{a}$, where the equation of state is allowed to vary. We adopt narrower, but still uninformative, flat priors on $H_{0}$ and $\Omega_{m}$. For $w_{0}$ and $w_{a}$, they are either fixed or allowed to vary within our uniform prior ranges. The priors are:
@@
For LVK we include in our analysis modes up to $\ell=400$, but the left plot of Fig.  11 shows that the information saturates at around $\ell\sim 200$. For the LVK+ET and ET+2CE networks the analysis includes multipoles up to $\ell\leq 700$, but the center and right plots of Fig.  11 shows that the constraint converges around $\ell=300$ in both cases. As for the redshift range, LVK already saturates for $z\lesssim 0.8$. However, the scenarios including ET may still benefit from galaxy or quasar catalogs that reach deeper than $z=2$ (which is the limit of the light cones we used in this work).
@@ 4.2 Consistency checks and possible systematics
The first test examines the impact of modifying the modeling of the matter power spectrum by disabling the small-scale corrections due to non-linear structure formation implemented by means of the HALOFIT model. We simply switch off these non-linear corrections in the theoretical model (which are then computed from simulations that use a linear matter power spectrum), while using as data the simulations made with the non-linear (HALOFIT) spectrum. From this analysis we have obtained the same results from our standard analysis (in which the non-linear corrections are included in the theoretical input) – see the left plot of Fig.  13. This highlights the insensitivity of the cross-correlation signal peak to non-linear corrections, demonstrating the robustness of the Peak Sirens method against different modeling of structure formation in the non-linear regime.
@@ 5 Discussion and conclusions
Our results show that, assuming $\Lambda$CDM and projected O5 sensitivity, we expect LVK should be able to measure $H_{0}$ with 4% precision by using an external dataset to fix $\Omega_{m}$ or with 7% precision by itself. This has the potential to provide a significant new outlook on the current $H_{0}$ tension. Looking further into the future, the upcoming third generation GW detectors should provide exquisite constraints on some key cosmological parameters with the cross-correlation technique. In particular a single such detector operating in tandem with LVK can constrain $H_{0}$ to 1% precision assuming $\Lambda$CDM, and to within 3% assuming a very flexible $w_{0}w_{a}$ model with uninformative priors. It can also constrain $\Omega_{m}$, $w_{0}$ and $w_{a}$ with very competitive precision.

### v2

@@ Abstract
Observations of gravitational waves (GWs) from dark sirens allow us to infer their locations and distances. Galaxies, on the other hand, have precise angular positions but no direct measurement of their distances – only redshifts. The cross-correlation of GWs, which we limit here to binary black hole mergers (BBH), in spherical shells of luminosity distance $D_{L}$, with galaxies in shells of redshift $z$, leads to a direct measurement of the Hubble diagram $D_{L}(z)$. Since this standard ruler relies only on the statistical proximity of the dark sirens and galaxies (a general property of large-scale structures), it is essentially model-independent: the correlation is maximal when both redshift and $D_{L}$ shells coincide. We forecast the constraining power of this technique, which we call Peak Sirens, for run 5 (O5) of LIGO-Virgo-KAGRA (LVK), as well as for the third-generation observatories Einstein Telescope and Cosmic Explorer. We employ thousands of full-sky light cone simulations with realistic numbers for the tracers, and include masking by the Milky Way, lensing and inhomogeneous GW sky coverage. We find that the method is not expected to suffer from some of the issues present in other dark siren methods, such as biased constraints due to incompleteness of galaxy catalogs or dependence on priors for the merger rates of BBH. We show that with Peak Sirens, given the projected O5 sensitivity, LVK can measure $H_{0}$ with $7\%$ precision by itself, assuming $\Lambda$CDM, and $4\%$ precision using external datasets to constrain $\Omega_{m}$. We also show that future third-generation GW detectors can achieve, without external data, sub-percent uncertainties in $H_{0}$ assuming $\Lambda$CDM, and 3% in a more flexible $w_{0}w_{a}$ model. The method also shows remarkable robustness against systematic effects such as the modeling of non-linear structure formation.
@@ 1 Introduction
The large-scale structures of the Universe are increasingly being mapped with multiple tracers and messengers, and the complementarity between different surveys has enormous potential to test cosmological models and resolve some of the existing tensions of the standard $\Lambda$CDM scenario [1, 2].
In this paper we investigate how we can exploit the cross-correlations between redshift surveys and detections of gravitational waves (GW) from dark sirens (most of which are binary black hole mergers, BBH) to trace the Hubble distance diagram $D_{L}(z)$ – even when the galaxies that host GWs are absent from the redshift survey. The idea is illustrated in Fig. 1: a galaxy map, where the radial information is only in the redshifts, covers the same area in the sky as a map of GWs, for which the gravitational wave (luminosity) distances can be inferred from the waveforms.
@@ 2.1 Multi-tracer likelihoods
2.2 A fast estimator for the theoretical redshift-distance cross-correlations
@@
[…]
Suppose that the luminosity distance bins of our data set are fixed – that is, independent of the cosmology. In practice, this is how we would proceed to compute the cross correlations of a set of observed BBHs with a set of galaxies in fixed redshift bins. Now let’s say that the small bins $D_{\lambda}$ are not fixed, but rather we fix some set of very thin redshift bins $z_{\lambda}$. As the connection between redshift and distance is mediated by the cosmological model, we can write $D_{\lambda}=D_{L}(z_{\lambda},\theta)\,$. Therefore, we can write the harmonic mode in the distance bin $D_{i}$ as a sum over the redshift bins $z_{\lambda}$ that fit into that distance bin, for the given cosmology:
[…]
The procedure above implements the calculation of the angular power spectra in practice, for the given set of parameters $\theta$. Now say that we change the set of cosmological parameters from some fiducial ${\theta}_{\rm fid}$ to another arbitrary set ${\theta}$. Although the amplitudes $a_{\ell m}^{b}(z_{\lambda},\theta)$ change, the correlation function still peaks when the redshift of the galaxies coincide with $z_{\lambda}$. Hence, the main effect of changing the cosmological parameters is the shift in the location of the peak in $D_{L}(z)$, which can be represented by a remapping of the redshift bins ($z_{\lambda}$) into the distance bins $D_{i}$ in Eq. (2.13).
Let us then approximate the set of harmonic coefficients in an arbitrary cosmology to those coefficients computed from the distribution of BBHs (and galaxies) in the fiducial model rescaled to the distances in the new cosmology,
$a_{\ell m}^{b}(D_{i},\theta)\,\equiv\sum_{z_{\lambda}\in\,z(D_{i},\theta)}\,a_{\ell\,m}^{b}(z_{\lambda},\theta_{\rm fid})\,.$ (2.14)
@@
where $C^{gb}_{\ell}(z_{j},z_{\lambda},\theta_{\rm fid})$ is the cross-angular power spectrum computed in thin redshift bins $z_{\lambda}$ (both for galaxies and BBHs) and in a fiducial cosmology $\theta_{\rm fid}$, averaged over a given number of realizations (for our purposes, of the order of thousands, as we will see). The only general condition is that the distance bins are wide enough to comprise several thin redshift bins, so that our analysis is not significantly affected by that spatial resolution.
The main reason why Eq. (2.15) is a good approximation derives from the coincidence of the peak of the cross-correlation with the Hubble diagram, which is a generic feature of the large-scale structures and as such should not depend significantly on the galaxy or BBH populations, bias modeling or non-linearities. The analytical argument is as follows: for a given set of cosmological parameters $\mathbf{{\theta}}$, the angular power spectrum for the cross-correlations between galaxies and BBHs is given by:
[…]
where $b_{g}$ and $b_{b}$ are the biases of the galaxies and BBHs, and $\chi$ is the comoving distance. In the Limber approximation, this turns into
$C_{\ell}(z_{g},D_{L},\mathbf{{\theta}})\approx\frac{\delta\left[\chi(z_{g},\mathbf{{\theta}})-\chi(D_{L},\mathbf{{\theta}})\right]}{\chi(z_{g},\mathbf{{\theta}})^{2}}\,b_{g}(z_{g})\,b_{b}(D_{L})\,P\left[k=\frac{\ell+1/2}{\chi(z_{g},\mathbf{{\theta}})},z_{g},\mathbf{\theta}\right]\,.$ (2.17)
In the equation above, only the cosmological parameters that affect distance (mainly, $H_{0}$ and $\Omega_{m}$) come into play inside the Dirac delta function, while the other parameters (as well as the biases) affect only the amplitude of the spectrum. Hence, the signal is dictated primarily by the sharp signature of the Dirac delta function, which enforces the match between $\chi(z_{g},\mathbf{\theta})$ and $\chi(D_{L},\mathbf{\theta})$.
Equation (2.15) is the key approximation that makes our analysis possible. As we have shown above, the log-likelihood (2.1) will be sensitive mostly to the peak of the cross-correlation signal, broadened by the uncertainty in the BBH localization, and this is overwhelmingly dictated by the distance-redshift relation. The validity of this approximation will be demonstrated in practice, using simulations, in Section 4.2.
@@ 3 Simulations
In order to assess the potential of the Peak Sirens method to constrain cosmology we forward-model the cosmological distribution of galaxies and BBHs. We will make use of two sets of simulations: the first will be used to compute the theoretical $C_{\ell}^{gb}$; the second will be used to forecast future data. This first set of simulations is a necessary part of the Peak Sirens method, and must always be performed. The second set is only used to make forecasts, and will be replaced by the real data when it becomes available. Clearly one does not know a priori which is the right cosmology for the first set, but we will show below that the method is largely insensitive to the choice of cosmology.
The starting point is a full-sky cosmological mock of the past light cone, possible thanks to the publicly available GLASS code [43]. We used GLASS to simulate 1000 lognormal galaxy mocks in very thin redshift bins, of width $dz=0.02$, over the interval $z\in[0,2]$ – resulting in the mean redshift bins $\bar{z}$:
@@ 3.1 Gravitational wave Fisher forecasts
We performed Fisher forecasts using the code GWDALI [63], [^3] Here we employ only the Fisher matrix and not the full DALI method with doublet or triplet [64]. with the following 11 free parameters:
@@
where $M_{c}$ is the redshifted chirp mass, $\eta$ is the symmetric mass ratio, and $t_{c}$ and $\phi_{c}$ are the time and phase of coalescence. We considered as threshold for detection having the total SNR of the network greater than 12. For the LVK+ET network, we have further imposed that at least one of the LVK detectors have SNR > 1. That does not mean, therefore, that the events included in this case would actually be detected by the 2nd generation network. Rather, LVK would help a 3rd generation detector (here, the triangular ET) to triangulate the signal improving localization knowledge, similarly to how Virgo has helped LIGO on localizing sources in their first three observing runs even if for the events of which Virgo alone can not claim a detection. After these cuts in SNR, we prospect the total amount of 377 observed events for LVK O5, while for LVK+ET and ET+2CE we expect this number to grow to, respectively, 54991 and 73260 detected events.
@@
Figure 3: Cumulative distribution of forecast sky position uncertainties (left) and relative distance errors (right) for the three GW detector networks here considered. The number of events is much higher for LVK+ET compared to LVK O5, especially at high $z$, a fraction of which with poor localization, result in a modest improvement in the average localization. See also Figure 4.
@@ 3.2 Galaxy and BBH mocks
with $z_{0}=0.7$, $\alpha=2.0$, $\beta=1.5$, and a linear galaxy bias, $b_{g}(z)=(1+z)$. This prescription yields a good fit for magnitude-limited optical surveys, as it captures in a simple formula volume effects at low redshifts, the evolution of the galaxy luminosity function, as well as decreasing counts due to increasingly fainter objects at high redshifts. With our choice of parameters (including the normalization of the Smail distribution) we obtained approximately 34 million galaxies up to $z=2$, a number that will easily be surpassed by upcoming surveys. To account for partial sky coverage, we apply a mask to the galaxy maps, restricting the analysis to latitudes $|\phi|\geq 19^{\circ}$, approximately $\nicefrac{{1}}{{3}}$ of the sky.
@@ 3.2.3 Weak lensing effects
Fig. 7 can also be directly compared with the analytical prediction of Ref. [4] – see Fig. 1 of that paper. Apart from an overall normalization, the results for the weak-lensing contribution to the signal of $C_{\ell}^{gb}$ for the optimal case of ET+2CE are consistent with that analytical estimate. However, due to the long tail on the convergence PDFs, we observe a non-negligible correlation between BBHs and galaxies at redshifts far below $z_{g}=1$. One of the main differences between our results and those of Ref. [4] is the steep drop of the lensing signal as $z\to 0$, which is due to the fact that the convergence PDFs become narrower and nearly Gaussian as we $z\to 0$, corresponding to a fast decreasing fraction of highly magnified objects.
@@ 4.1 Forecasts for cosmological parameters
The distance bins are kept fixed but can be defined arbitrarily, independently of the redshift bins. For this analysis we choose 50 distance bins ranging from $D_{\text{fid}}(z=0)$ to $D_{\text{fid}}(z=2)$, in equally spaced redshift intervals, where by $D_{\text{fid}}$ we mean the luminosity distance calculated in the fiducial cosmology. This binning strategy offers the practical advantage that it ensures that each distance bin includes one or more thin redshift bins in the discrete sum of Eq. (2.15).
@@
For middle- and long-term scenarios, we investigate LVK+ET and ET+2CE as potential networks of GW detectors. For these forecasts we also consider, besides $\Lambda$CDM, a second model, $w_{0}w_{a}$, where the equation of state is allowed to vary. We adopt narrower, but still uninformative, flat priors on $H_{0}$ and $\Omega_{m}$. For $w_{0}$ and $w_{a}$, they are either fixed or allowed to vary within our uniform prior ranges. The priors are:
@@
For LVK we include in our analysis modes up to $\ell=400$, but the left plot of Fig. 11 shows that the information saturates at around $\ell\sim 200$. For the LVK+ET and ET+2CE networks the analysis includes multipoles up to $\ell\leq 700$, but the center and right plots of Fig. 11 shows that the constraint converges around $\ell=300$ in both cases. As for the redshift range, LVK already saturates for $z\lesssim 0.8$. However, the scenarios including ET may still benefit from galaxy or quasar catalogs that reach deeper than $z=2$ (which is the limit of the light cones we used in this work).
@@ 4.2 Consistency checks and possible systematics
The first test examines the impact of modifying the modeling of the matter power spectrum by disabling the small-scale corrections due to non-linear structure formation implemented by means of the HALOFIT model. We simply switch off these non-linear corrections in the theoretical model (which are then computed from simulations that use a linear matter power spectrum), while using as data the simulations made with the non-linear (HALOFIT) spectrum. From this analysis we have obtained the same results from our standard analysis (in which the non-linear corrections are included in the theoretical input) – see the left plot of Fig. 13. This highlights the insensitivity of the cross-correlation signal peak to non-linear corrections, demonstrating the robustness of the Peak Sirens method against different modeling of structure formation in the non-linear regime.
@@ 5 Discussion and conclusions
Our results show that, assuming $\Lambda$CDM and projected O5 sensitivity, we expect LVK should be able to measure $H_{0}$ with 4% precision by using an external dataset to fix $\Omega_{m}$ or with 7% precision by itself. This has the potential to provide a significant new outlook on the current $H_{0}$ tension. Looking further into the future, the upcoming third generation GW detectors should provide exquisite constraints on some key cosmological parameters with the cross-correlation technique. In particular, a single such detector operating in tandem with LVK can constrain $H_{0}$ to 1% precision assuming $\Lambda$CDM, and to within 3% assuming a very flexible $w_{0}w_{a}$ model with uninformative priors. It can also constrain $\Omega_{m}$, $w_{0}$ and $w_{a}$ with very competitive precision.